How much heat will be absorbed by a 72.3 g piece of aluminum (specific heat = 0.930 J/g・°C) as it changes temperature from 23.0°C to 67.0°C?
Answer #1

The absorbed heat is calculated:

q = m * sh * ΔT = 72.3 g * 0.930 J / g * ° C * (67 - 23) ° C = 2958.5 J

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.